Problem 10.40

A soccer ball, which has a circumference of 70.0 cm, rolls 12.0 yards in 3.45 s .

Part A

What was the average angular speed of the ball during this time? (units= rad/s)
Answer #1

circumference, P = 70 cm

r = P/(2*pi)

r = radius of ball = 0.7/(2*pi) = 0.111 m

distance traveled = 12 yard = 10.97 m

time taken = 3.45 sec

v = speed = distance travelled/time

v = 10.97/3.45 = 3.18 m/sec

Now we know that

V = w*r

w = V/r = 3.18/0.111 = 28.65 rad/sec